OpenStudy (anonymous):

The perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm². What are the dimensions of the rectangle? 11 cm by 16 cm 8 cm by 22 cm 5.5 cm by 32 cm 4 cm by 44 cm

hero (hero):

I'm going with the first choice

hero (hero):

P = 2(x + y) = 2(11 + 16) = 2(27) = 54 A = xy = 11*16 = 176